# Beyond Multi-Token Prediction: Pretraining LLMs with Future Summaries

### v1

@@ Abstract
Next-token prediction (NTP) has driven the success of large language models (LLMs), but it struggles with long-horizon reasoning, planning, and creative writing, with these limitations largely attributed to teacher-forced training. Multi-token prediction (MTP) partially mitigates these issues by predicting several future tokens at once, but it mostly captures short-range dependencies and offers limited improvement. We propose future summary prediction (FSP), which trains an auxiliary head to predict a compact representation of the long-term future, preserving information relevant for long-form generations. We explore two variants of FSP: handcrafted summaries, for example, a bag of words summary of the future of the sequence, and learned summaries, which use embeddings produced by a reverse language model trained from right to left. Large-scale pretraining experiments (3B and 8B-parameter models) demonstrate that FSP provides improvements over both NTP and MTP across math, reasoning, and coding benchmarks.
@@ 1 Introduction
Early progress in large language models was primarily driven by massive scaling of both data and compute (Brown et al., 2020; Kaplan et al., 2020). However, the returns from this vanilla scaling approach are beginning to diminish as we encounter the “data wall” (Sutskever, 2024). This has renewed efforts toward algorithmic advances, including new architectures and pretraining objectives, that can extract more predictive signal from a fixed amount of training data.
@@
Next-token prediction (NTP) with teacher forcing—training models by conditioning on ground-truth history when predicting the next token—is foundational to current pretraining methods. However, this approach introduces a train-inference mismatch known as exposure bias: during inference, the model must rely on its own outputs rather than the ground truth, leading to compounding errors and degraded long-range generation quality (Bengio et al., 2015). Moreover, teacher forcing can induce training-time shortcut learning as well, where the model exploits local cues from the ground-truth prefix instead of capturing true long-range dependencies (Bachmann and Nagarajan, 2024). These issues manifest most clearly in tasks demanding extended reasoning, narrative coherence, and open-ended creativity (Papalampidi et al., 2022; Nagarajan et al., 2025).
@@
An appealing alternative to teacher forcing is teacherless training, where the model learns from its own generations rather than relying on ground-truth histories. However, this approach is computationally intensive and challenging to parallelize. As a practical compromise, recent research has explored multi-token prediction (MTP) methods (Gloeckle et al., 2024), which train auxiliary heads to predict several future tokens simultaneously. MTP has demonstrated success in large-scale systems such as DeepSeek-V3 (Liu et al., 2024) and Qwen-3 (Yang et al., 2025). In MTP, each time step augments the standard next-token prediction ($x_{t+1}$) with auxiliary heads that predict additional future tokens (e.g., $x_{t+2}$ and beyond). However, these methods typically assume independence among the predicted tokens given the prefix, resulting in poor approximations of the true joint distribution over long future spans.
@@
In this work, we propose a different approach: rather than predicting multiple future tokens individually, we train a single auxiliary head to predict a summary representation of the future sequence. It aims to push as much information as possible about the future into a single target vector, while filtering out information that is inherently unpredictable and would only introduce noise. At each time step $t$, given the future tokens $(x_{t+2},x_{t+3},\dots,x_{t+\tau})$, we construct a summary vector $a(t,\tau)$ to provide supervision to the auxiliary head. We explore two complementary approaches for constructing future summaries: a simple token-level aggregation method, and our main contribution, a learned representation of the future sequence that captures long-range dependencies more effectively.
@@
  • Learned summary. Handcrafted summaries such as the one proposed above can be noisy: not all future tokens are equally relevant, and predicting them all can dilute the signal. To address this, we propose to learn a compact summary of the future. We do this by training a reverse language model ($g_{r}$) on reversed sequences, so that its hidden representation $a(t,\tau)=g_{r}(x_{\geq t+2})$ serves as a rich embedding of the future context. The auxiliary head of the forward model is then trained to match this representation with an $\ell_{2}$ loss.
Figure 1: A comparison of future-aware pretraining objectives. All methods take a prefix $x_{\leq t}$ as input. NTP: Only predicts the immediate next token. MTP: Uses multiple auxiliary heads, each predicting a specific future token. FSP-BCE: Our proposed hand-crafted summary method that predicts a multi-hot “bag-of-tokens” summary of a long future window using a single auxiliary head. FSP-RevLM: Our proposed learned summary method predicts a compact hidden representation of the future, which is generated by a reverse language model (RevLM).
@@ 2.1 Background on Next-Token and Multi-Token Prediction
Let us define a sequence of tokens as $X=(X_{1},X_{2},\dots,X_{T})$ sampled from a distribution $\mathbb{P}_{X}$, where each token is a discrete random variable supported on $\{1,2,\dots,V\}$.
@@
We discuss other variants of MTP such as DeepSeek-MTP and random future token MTP (Thankaraj et al., 2025) in Appendix 7.
@@ 3 Analyzing Future Summaries
To better understand how FSP mitigates the challenges of long-horizon planning, we consider the graph modeling benchmarks introduced in prior works. Our analysis yields following insights:
@@ 3.1 Long future summary is important
Figure 3: Analysis of FSP-BCE on the path-star task, which tests long-horizon planning. Left: Accuracy (mean over $5$ random seeds) of different pretraining objectives on degree 22 graphs with path lengths 66 and 88. Standard NTP generalizes poorly, and MTP’s accuracy degrades as the path length increases, while FSP-BCE achieves perfect accuracy. Right: An illustration of why NTP fails while FSP-BCE succeeds, where the input context is shown in grey and the target is in beige.
We consider the path-star graph, a directed acyclic graph (DAG) $G(d,l)$ composed of $d$ paths, each of length $l$, originating from a central start node $v_{\text{start}}$. The model is provided with the adjacency list of the graph in the prefix, and the task is to generate the path from $v_{\text{start}}$ to a designated end node $v_{\text{end}}$.
@@
A natural remedy is to reduce teacher forcing via future prediction approaches, which require the model to predict tokens further ahead. This makes shortcuts less effective, as multiple lookups in the adjacency list would be needed to predict future tokens. Hence, we consider the MTP approach, which predicts the immediate future token, and the handcrafted summary method FSP-BCE, which compresses the information about all the future tokens from the path. By summarizing the entire future trajectory, FSP-BCE substantially reduces teacher forcing, encouraging the model to plan the full path instead of exploiting shortcuts.
@@
We conduct experiments for two graphs $G(2,6)$ and $G(2,8)$, with all the approaches pretrained from scratch using the GPT-Mini architecture (details in Section˜8.1). At inference, we discard the auxiliary head used in future prediction, and the task is to generate the complete path given the prefix $p$. The evaluation metric checks whether the generated path exactly matches the true path. Results in Figure˜3 (left) show that both NTP and MTP fail to generalize (both obtained perfect training accuracy), and the accuracy of MTP degrades further for the scenario with longer path $G(2,8)$. Hence, predicting just the immediate future token is not enough, and FSP-BCE tackles this by efficiently compressing all the future tokens from the path, enabling it to achieve perfect accuracy in both cases. Appendix 8.1 (Table 4) shows that increasing the number of auxiliary heads in MTP can provide some improvement, but practical limits exist: even with four additional future heads MTP cannot solve the longer path graph $G(2,8)$.
@@ 3.2 Adaptive future summary is important
Figure 4: Analysis on the modified sibling discovery task which requires adaptive future summaries. Left: Convergence speed (mean (s.e.) over $3$ random seeds) relative to NTP, where lower values imply faster convergence. FSP-RevLM converges faster than NTP while FSP-BCE improves only in the cases with few components. Right: Task setup illustration– given the prefix, only future tokens in the highlighted component are informative. FSP-BCE, which summarizes all future tokens, suffers from irrelevant information, whereas FSP-RevLM summarizes only the informative aspects.
@@
We verify this empirically by comparing FSP-BCE and FSP-RevLM in experiments with varying number of total components. All models are pretrained from scratch with the GPT-Mini architecture (details in Section˜8.2), and the inference task requires generating a coherent sequence (siblings followed by parent for each component). At convergence, all models produce coherent sequences; to quantify the learning speedup over NTP, we report the ratio of steps to convergence relative to NTP, with lower values indicating faster learning. Results in figure 4 (left) shows that FSP-BCE improves over NTP only when the number of components is small, with gains disappearing for more than six components. In contrast, FSP-RevLM consistently achieves faster convergence across all component sizes, confirming the benefit of adaptive future summaries.
@@ 4.1 Setup
We pretrain 3B- and 8B-parameter models on corpora of 250B and 1T tokens, respectively, covering diverse domains. The majority of the data comes from DCLM-like sources and GitHub repositories, supplemented by specialized material in mathematics, programming, and related areas. Models are evaluated across a diverse set of benchmarks: ARC-Easy/Challenge (Clark et al., 2018) for general reasoning, MBPP (Austin et al., 2021) and HumanEval+ (Liu et al., 2023) for code generation, and GSM8K (Cobbe et al., 2021) and Math-500 (Hendrycks et al., 2021) for mathematical reasoning. Details of the pretraining corpus and hyperparameters are provided in Section˜8.3.
We first benchmark our primary method, FSP-RevLM, against the baselines: next-token prediction (NTP), standard multi-token prediction (MTP) (Gloeckle et al., 2024), and DeepSeek-MTP (DS-MTP) (Liu et al., 2024). A natural way to improve MTP for longer horizons is to add multiple auxiliary heads, each predicting a token farther into the future, but this approach quickly becomes impractical. We therefore constrain both MTP and DS-MTP to a single auxiliary head predicting the immediate future token. This design choice keeps the comparison consistent and aligned with the proposed FSP, since it uses a single auxiliary head.
Building on this unified single auxiliary head framework, we conduct an analysis of how MTP can be enhanced by predicting richer future targets instead of the immediate future token. In addition to predicting the learned future summary (FSP-RevLM), we compare handcrafted future summaries over short- and long-range windows. This includes the proposed multi-hot future summary (FSP-BCE) as a contributed baseline, and a random-token summary baseline (radomly sampling token from future), in line with prior works (Thankaraj et al., 2025; Gerontopoulos et al., 2025).
Note. All experiments are conducted under iso-data conditions, meaning that all the methods are trained on identical datasets. For the proposed FSP-RevLM, this implies that both the forward and reverse models are trained on the same data. In line with standard practice in distillation, we do not perform iso-compute comparisons that include the teacher model’s (ReverseLM) cost in the reported compute budget. In practical scenarios, the computational cost of training the teacher model is typically amortized, hence it can be treated as a one-time overhead that is excluded from comparisons of student models (Gemma et al., 2024, 2025).
@@ 4.2 Results
At the 3B scale (Table 2), DeepSeek-MTP is a strong baseline and obtains better performance than FSP-RevLM on most tasks, except math reasoning. More importantly, FSP-RevLM exhibits larger relative improvements as scale increases from 3B to 8B parameters, and becomes more favourable than DS-MTP. Further, note that even at the 3B scale, FSP-RevLM still beats MTP on ARC and math reasoning tasks, and performs comparably on MBPP, suggesting that even at smaller scales, learning to predict future summaries may provide more effective auxiliary signal than immediate future token prediction with MTP.
@@
Table 1: Pretraining at 8B scale. We benchmark the proposed FSP-RevLM approach against NTP, MTP, and DS-MTP. Results (mean $\pm$ s.e. over 3 seeds) report pass@16 for code/math tasks and accuracy for ARC. FSP-RevLM achieves the strongest overall performance, with large gains over the baselines on ARC tasks and MATH, and competitive results with (DS) MTP on code benchmarks.
@@
Table 2: Pretraining at 3B scale. We benchmark the proposed FSP-RevLM approach against NTP, MTP, and DS-MTP. Results (mean $\pm$ s.e. over 3 seeds) report pass@16 for code/math tasks and accuracy for ARC. At this smaller scale, DS-MTP is a strong overall baseline, but FSP-RevLM outperforms it on math reasoning tasks, and also provides substantial gains over MTP on ARC and math reasoning tasks. Further, as we scale the approaches to 8B parameters, FSP-RevLM scales more favorably than DS-MTP, overtaking it on most tasks.
4.3 Analyzing MTP with different future summaries
@@
Our results show that random-token handcrafted future summaries (MTP-Skip) perform worse than standard MTP with immediate future token prediction, and performance further degrades as the future window ($\tau$) increases. In contrast, the proposed multi-hot or bag-of-words handcrafted future summaries yield meaningful improvements over MTP, especially on math reasoning tasks, with both shorter $(\tau=12)$ and longer $(\tau=100)$ future windows. For example, FSP-BCE with $\tau=12$ achieves $33.1\%$ on MATH (+2.2 points) and $69.9\%$ on GSM8K (+2.1 points), while even longer windows ($\tau=100$) further amplifies the performance on GSM8K ($71.4\%$, +3.6 points).
@@
Finally, our learned future summaries (FSP-RevLM) outperform MTP across all evaluation tasks, with especially pronounced gains on math reasoning: $35.1\%$ on MATH (+4.2 points) and $70.5\%$ on GSM8K (+3.5 points). Further analysis (Figure 5) for these math reasoning tasks shows that learned summaries promote greater output diversity across different pass@k settings, compared to vanilla MTP with immediate future token prediction.
@@
In Section˜8.3 (Table 5), we replicate these findings at 3B scale, where both handcrafted and learned future summaries improve over vanilla MTP, again most prominently on math reasoning tasks. Additional ablations explore the effects of omitting reweighting in FSP-BCE and predicting learned summaries from deeper layers of the reverseLM (FSP-RevLM).
Method MBPP GSM8K MATH HumanEval+ ARC-Challenge ARC-Easy MTP $0.672$ $(0.008)$ $0.678$ $(0.007)$ $0.309$ $(0.006)$ $\bm{0.541}$ $(0.011)$ $0.552$ $(0.000)$ $0.736$ $(0.000)$ MTP-Skip $\tau$:4 $0.658$ $(0.005)$ $0.639$ $(0.004)$ $0.277$ $(0.020)$ $0.508$ $(0.009)$ $0.494$ $(0.003)$ $0.722$ $(0.000)$ MTP-Skip $\tau$:12 $0.623$ $(0.002)$ $0.621$ $(0.005)$ $0.287$ $(0.018)$ $0.486$ $(0.010)$ $0.512$ $(0.000)$ $0.710$ $(0.003)$ MTP-Skip $\tau$:32 $0.611$ $(0.008)$ $0.598$ $(0.007)$ $0.271$ $(0.005)$ $0.459$ $(0.007)$ $0.379$ $(0.000)$ $0.564$ $(0.000)$ FSP-BCE $\tau$:12 $0.669$ $(0.005)$ $0.699$ $(0.006)$ $0.331$ $(0.016)$ $0.508$ $(0.005)$ $\bm{0.562}$ $\bm{(0.000)}$ $0.737$ $(0.000)$ FSP-BCE $\tau$:100 $0.671$ $(0.002)$ $\bm{0.714}$ $\bm{(0.009)}$ $0.331$ $(0.007)$ $0.500$ $(0.019)$ $0.459$ $(0.000)$ $0.662$ $(0.000)$ FSP-RevLM $0.683$ $(0.006)$ $0.705$ $(0.004)$ $\bm{0.351}$ $\bm{(0.017)}$ $\bm{0.541}$ $(0.009)$ $0.559$ $(0.000)$ $\bm{0.766}$ $\bm{(0.000)}$
Table 3: Analysis of future-summary strategies at 8B Scale. We evaluate the effect of different future-summary prediction approaches against vanilla MTP. Results (mean $\pm$ s.e. over 3 seeds) report pass@16 for code/math tasks and accuracy for ARC. Handcrafted random-token future summaries (MTP-Skip) perform worse than standard MTP. In contrast, handcrafted multi-hot summaries (FSP-BCE) improve over standard MTP, especially on math reasoning (e.g., GSM8K and MATH), while learned summaries (FSP-RevLM) provide the largest gains across math reasoning and ARC tasks.

### v2

@@ Abstract
Next-token prediction (NTP) has driven the success of large language models (LLMs), but it struggles with long-horizon reasoning, planning, and creative writing, with these limitations largely attributed to teacher-forced training. Multi-token prediction (MTP) partially mitigates these issues by predicting several future tokens at once, but it mostly captures short-range dependencies and offers limited improvement. We propose future summary prediction (FSP), which trains an auxiliary head to predict a compact representation of the long-term future, preserving information relevant for long-form generations. We explore two variants of FSP: handcrafted summaries, for example, a bag of words summary of the future sequence, and learned summaries, which use embeddings produced by a reverse language model trained from right-to-left order. Large-scale pretraining experiments (3B and 8B-parameter models) demonstrate that FSP provides improvements over both NTP and MTP across math, reasoning, and coding benchmarks.
@@ 1 Introduction
Early progress in large language models was primarily driven by massive scaling of both data and compute (Brown et al., 2020; Kaplan et al., 2020). However, the returns from this vanilla scaling approach are beginning to diminish as we encounter the “data wall” (Villalobos et al., 2024; Sutskever, 2024). This has renewed efforts toward algorithmic advances, including new architectures and pretraining objectives, that can extract more predictive signal from a fixed amount of training data (Muennighoff et al., 2023).
@@
Next-token prediction (NTP) with teacher forcing—training models by conditioning on ground-truth history when predicting the next token—is foundational to current pretraining methods. However, this approach introduces a train-inference mismatch known as exposure bias: during inference, the model must rely on its own outputs rather than the ground truth, leading to compounding errors and degraded long-range generation quality (Bengio et al., 2015; Ranzato et al., 2015). Moreover, teacher forcing can induce training-time shortcut learning as well, where the model exploits local cues from the ground-truth prefix instead of capturing true long-range dependencies (Bachmann and Nagarajan, 2024). These issues manifest most clearly in tasks demanding long-horizon planning, narrative coherence, and open-ended creativity (Papalampidi et al., 2022; Nagarajan et al., 2025).
@@
An appealing alternative to teacher forcing is teacherless training, where the model learns from its own generations rather than relying on ground-truth histories. However, this approach is computationally intensive and challenging to parallelize. As a practical compromise, recent research has explored multi-token prediction (MTP) methods (Gloeckle et al., 2024), which train auxiliary heads to predict several future tokens simultaneously. MTP has demonstrated success in large-scale systems such as DeepSeek-V3 (Liu et al., 2024), Qwen-3 (Yang et al., 2025), MiMo (Xiaomi et al., 2025), and Nemotron-3 (Blakeman et al., 2025). In MTP, each time step augments the standard next-token prediction ($x_{t+1}$) with auxiliary heads that predict additional future tokens (e.g., $x_{t+2}$ and beyond). However, these methods typically assume conditional independence among the predicted tokens given the prefix, resulting in poor approximations of the true joint distribution over long future spans.
@@
In this work, we propose a different approach: rather than predicting multiple future tokens individually, we train a single auxiliary head to predict a summary of the future sequence. It aims to push as much information as possible about the future into a single target vector, while filtering out information that is inherently unpredictable and would only introduce noise. At each time step $t$, given the future tokens $(x_{t+2},x_{t+3},\dots,x_{t+\tau})$, we construct a summary vector $a(t,\tau)$ to provide supervision to the auxiliary head. We explore two complementary approaches for constructing future summaries: a simple token-level aggregation method, and our main contribution, a learned representation of the future sequence that captures long-range dependencies more effectively.
@@
  • Learned summary. Handcrafted summaries such as the one proposed above can be noisy as not all future tokens are equally relevant. To address this, we propose to learn a compact summary of the future. We do this by training a reverse language model ($g_{r}$) on reversed sequences (right-to-left order), so that its hidden representation $a(t,\tau)=g_{r}(x_{\geq t+2})$ serves as a rich embedding of the future context. The auxiliary head of the forward model is then trained to match this representation with an $\ell_{2}$ loss.
Figure 1: A comparison of future-aware pretraining objectives. All methods take a prefix $x_{\leq t}$ as input. NTP: Only predicts the immediate next token. MTP: Uses multiple auxiliary heads, each predicting a specific future token. FSP-BoW: Our proposed hand-crafted summary method that predicts a multi-hot “bag-of-tokens” summary of a long future window using a single auxiliary head. FSP-RevLM: Our proposed learned summary method predicts a compact hidden representation of the future, which is generated by a reverse language model (RevLM).
@@ 2.1 Background on Next-token and Multi-token Prediction
Let define a sequence of tokens as $X=(X_{1},X_{2},\dots,X_{T})$ sampled from a distribution $\mathbb{P}_{X}$, where each token is a discrete random variable supported on $\{1,2,\dots,V\}$.
@@
We discuss other variants of MTP such as DeepSeek-MTP (Liu et al., 2024), Joint-token Prediction (Ahn et al., 2025), and random future token MTP (Thankaraj et al., 2025) in Appendix 8.
@@ 3 Analyzing Future Summaries
Future summary prediction provides a unified view of a broad family of pretraining objectives, covering MTP and its variants that sample random future tokens (Thankaraj et al., 2025; Gerontopoulos et al., 2025), and our proposed bag-of-words (FSP-BoW) and ReverseLM (FSP-RevLM) summary prediction objectives.
To better understand how the FSP framework clarifies the tradeoffs of each objective on long-horizon planning, we consider the graph modeling benchmarks introduced in prior works. Our analysis yields following insights:
@@ 3.1 Long future summary is important
Figure 3: Analysis of FSP-BoW on the path-star task, which tests long-horizon planning. Left: Accuracy (mean over 55 random seeds) of different pretraining objectives on degree 22 graphs with path lengths 66 and 88. Standard NTP generalizes poorly, and MTP’s accuracy degrades as the path length increases, while FSP-BoW achieves perfect accuracy. Right: An illustration of why NTP fails while FSP-BoW succeeds, where the input context is shown in grey and the target is in beige.
We consider the path-star graph, a directed acyclic graph (DAG) $G(d,l)$ composed of $d$ paths, each of length $l$, originating from a central start node $v_{\text{start}}$. The model is given the adjacency list of the graph in the prefix, and the task is to generate the path from $v_{\text{start}}$ to a designated end node $v_{\text{end}}$.
@@
A natural remedy is to reduce teacher forcing via future prediction approaches, which require the model to predict tokens further ahead. This makes shortcuts less effective, as multiple lookups in the adjacency list would be needed to predict future tokens. Hence, we consider the MTP approach, which predicts the immediate future token, and the handcrafted summary method FSP-BoW, which compresses the information about all the future tokens from the path. By summarizing the entire future trajectory, FSP-BoW substantially reduces teacher forcing, encouraging the model to plan the full path instead of exploiting shortcuts.
@@
We conduct experiments for two graphs $G(2,6)$ and $G(2,8)$, with all the approaches pretrained from scratch using the GPT-Mini architecture (details in Section˜9.1). At inference, we discard the auxiliary head used in future prediction, and the task is to generate the complete path given the prefix $p$. The evaluation metric checks whether the generated path exactly matches the true path. Results in Figure˜3 (left) show that both NTP and MTP fail to generalize (both obtained perfect training accuracy), and the accuracy of MTP degrades further for the scenario with longer path $G(2,8)$. Hence, predicting just the immediate future token is not enough, and FSP-BoW tackles this by efficiently compressing all the future tokens from the path, enabling it to achieve perfect accuracy in both cases. Appendix 9.1 (Table 4) shows that increasing the number of auxiliary heads in MTP can provide some improvement, but practical limits exist: even with four additional future heads MTP cannot solve the longer path graph $G(2,8)$.
@@ 3.2 Adaptive future summary is important
Figure 4: Analysis on the modified sibling discovery task which requires adaptive future summaries. Left: Convergence speed (mean (s.e.) over 33 random seeds) relative to NTP, where lower values imply faster convergence. FSP-RevLM converges faster than NTP while FSP-BoW improves only in the cases with few components. Right: Task setup illustration– given the prefix, only future tokens in the highlighted component are informative. FSP-BoW, which summarizes all future tokens, suffers from irrelevant information, whereas FSP-RevLM summarizes only the informative aspects.
@@
We verify this empirically by comparing FSP-BoW and FSP-RevLM in experiments with varying number of total components. All models are pretrained from scratch with the GPT-Mini architecture (details in Section˜9.2), and the inference task requires generating a coherent sequence (siblings followed by parent for each component). At convergence, all models produce coherent sequences; to quantify the learning speedup over NTP, we report the ratio of steps to convergence relative to NTP, with lower values indicating faster learning. Since all methods eventually reach perfect consistency at convergence, time to convergence is a reliable metric for comparison. Results in figure 4 (left) shows that FSP-BoW improves over NTP only when the number of components is small, with gains disappearing for more than six components. In contrast, FSP-RevLM consistently achieves faster convergence across all component sizes, confirming the benefit of adaptive future summaries.
@@ 4.1 Setup
We pretrain 3B- and 8B-parameter models on corpora of 250B and 1T tokens, respectively, covering diverse domains. The majority of the data comes from DCLM-like sources and GitHub repositories, supplemented by specialized material in mathematics, programming, and related areas. Models are evaluated across a diverse set of benchmarks: ARC-Easy/Challenge (Clark et al., 2018) for general reasoning, MBPP (Austin et al., 2021) and HumanEval+ (Liu et al., 2023) for code generation, and GSM8K (Cobbe et al., 2021) and Math-500 (Hendrycks et al., 2021) for mathematical reasoning. Details of the pretraining corpus and hyperparameters are provided in Section˜10.1.
We first benchmark our primary method, FSP-RevLM, against the baselines: next-token prediction (NTP), multi-token prediction (MTP) (Gloeckle et al., 2024), and DeepSeek-MTP (DS-MTP) (Liu et al., 2024). Further, a natural way to improve MTP for longer horizons is to add multiple auxiliary heads, each predicting a token farther into the future, but this approach quickly becomes impractical. We therefore constrain both MTP and DS-MTP to a single auxiliary head predicting the immediate future token. This design choice keeps the comparison consistent and aligned with the proposed FSP, since it uses a single auxiliary head.
Building on this unified single auxiliary head framework, we conduct an analysis of how MTP can be enhanced by predicting richer future targets instead of the immediate future token. In addition to predicting the learned future summary (FSP-RevLM), we compare handcrafted future summaries over short- and long-range windows. This includes the proposed bag-of-words future summary (FSP-BoW) as a contributed baseline, and randomly sampled future tokens, in line with prior works (Thankaraj et al., 2025; Gerontopoulos et al., 2025).
Note regarding experiment design. All experiments are conducted under iso-data conditions, meaning that all the methods are trained on identical datasets. For the proposed FSP-RevLM, this implies that both the forward and reverse models are trained on the same data. In line with standard practice in distillation, we do not perform iso-compute comparisons that include the teacher model’s (ReverseLM) cost in the reported compute budget. In practical scenarios, the computational cost of training the teacher model is typically amortized, hence it can be treated as a one-time overhead that is excluded from comparisons of student models (Gemma et al., 2024, 2025).
Note regarding FSP-RevLM implementation. In our experiments with FSP-RevLM, the reverse model is the same size as the forward model (and other baselines), and it is trained for the same number of steps. As a result, FSP-RevLM roughly doubles the total compute cost compared to standard NTP training. While FSP-RevLM increases training FLOPs, we believe this tradeoff is reasonable in today’s compute-rich, data-limited scaling regime (Kim et al., 2025). The field has effectively hit the data wall, whereas available compute continues to grow. Progress increasingly depends on using this growing compute-per-token budget to extract more value from fixed datasets. In this context, methods that deliver measurable gains without requiring additional data, even at higher compute cost, are valuable.
@@ 4.2 Results
At the 3B scale (Table 2), DeepSeek-MTP is a strong baseline and obtains better performance than FSP-RevLM on most tasks, except math reasoning. More importantly, FSP-RevLM exhibits larger relative improvements as scale increases from 3B to 8B parameters, and becomes more favourable than DS-MTP. Note that even at the 3B scale, FSP-RevLM still beats MTP on ARC and math reasoning tasks, and performs comparably on MBPP, suggesting that even at smaller scales, learning to predict future summaries may provide more effective auxiliary signal than immediate future token prediction with MTP.
Finally, as a sanity check, in Appendix 10.2 we evaluate these models on standard language modeling task. We find that future-prediction objectives remain competitive with NTP for language modeling, indicating that the reasoning gains over NTP do not come at the cost of degraded language modeling performance.
@@
Table 1: Pretraining at 8B scale. We benchmark the proposed FSP-RevLM approach against NTP, MTP, and DS-MTP. Results (mean $\pm$ s.e. over 3 seeds) report pass@16 for code/math tasks and pass@1 for ARC. FSP-RevLM achieves the strongest overall performance, with large gains over the baselines on ARC tasks and MATH, and competitive results with (DS) MTP on code benchmarks.
@@
Table 2: Pretraining at 3B scale. We benchmark the proposed FSP-RevLM approach against NTP, MTP, and DS-MTP. Results (mean $\pm$ s.e. over 3 seeds) report pass@16 for code/math tasks and pass@1 for ARC. At this smaller scale, DS-MTP is a strong overall baseline, but FSP-RevLM outperforms it on math reasoning tasks, and also provides substantial gains over MTP on ARC and math reasoning tasks. Further, as we scale the approaches to 8B parameters, FSP-RevLM scales more favorably than DS-MTP, overtaking it on most tasks.
Method MBPP GSM8K MATH HumanEval+ ARC-Challenge ARC-Easy MTP $0.672$ $(0.008)$ $0.678$ $(0.007)$ $0.309$ $(0.006)$ $\bm{0.541}$ $(0.011)$ $0.552$ $(0.000)$ $0.736$ $(0.000)$ MTP-Skip $\tau$:4 $0.658$ $(0.005)$ $0.639$ $(0.004)$ $0.277$ $(0.020)$ $0.508$ $(0.009)$ $0.494$ $(0.003)$ $0.722$ $(0.000)$ MTP-Skip $\tau$:12 $0.623$ $(0.002)$ $0.621$ $(0.005)$ $0.287$ $(0.018)$ $0.486$ $(0.010)$ $0.512$ $(0.000)$ $0.710$ $(0.003)$ MTP-Skip $\tau$:32 $0.611$ $(0.008)$ $0.598$ $(0.007)$ $0.271$ $(0.005)$ $0.459$ $(0.007)$ $0.379$ $(0.000)$ $0.564$ $(0.000)$ FSP-BoW $\tau$:12 $0.669$ $(0.005)$ $0.699$ $(0.006)$ $0.331$ $(0.016)$ $0.508$ $(0.005)$ $\bm{0.562}$ $\bm{(0.000)}$ $0.737$ $(0.000)$ FSP-BoW $\tau$:100 $0.671$ $(0.002)$ $\bm{0.714}$ $\bm{(0.009)}$ $0.331$ $(0.007)$ $0.500$ $(0.019)$ $0.459$ $(0.000)$ $0.662$ $(0.000)$ FSP-RevLM $0.683$ $(0.006)$ $0.705$ $(0.004)$ $\bm{0.351}$ $\bm{(0.017)}$ $\bm{0.541}$ $(0.009)$ $0.559$ $(0.000)$ $\bm{0.766}$ $\bm{(0.000)}$
Table 3: Analysis of future summary strategies at 8B Scale. We evaluate the effect of different future-summary prediction approaches against vanilla MTP. Results (mean $\pm$ s.e. over 3 seeds) report pass@16 for code/math tasks and pass@1 for ARC. Handcrafted random-token future summaries (MTP-Skip) perform worse than standard MTP. In contrast, handcrafted multi-hot summaries (FSP-BoW) improve over standard MTP, especially on math reasoning (e.g., GSM8K and MATH), while learned summaries (FSP-RevLM) provide the largest gains across math reasoning and ARC tasks.
Analyzing MTP with different future summaries.
@@
Our results show that random-token handcrafted future summaries (MTP-Skip) perform worse than standard MTP with immediate future token prediction, and performance further degrades as the future window ($\tau$) increases. In contrast, the proposed multi-hot or bag-of-words handcrafted future summaries yield meaningful improvements over MTP, especially on math reasoning tasks, with both shorter $(\tau=12)$ and longer $(\tau=100)$ future windows. For example, FSP-BoW with $\tau=12$ achieves $33.1\%$ on MATH (+2.2 points) and $69.9\%$ on GSM8K (+2.1 points), while even longer windows ($\tau=100$) further amplifies the performance on GSM8K ($71.4\%$, +3.6 points).
@@
Finally, our learned future summaries (FSP-RevLM) outperform MTP across all evaluation tasks, with especially pronounced gains on math reasoning: $35.1\%$ on MATH (+4.2 points) and $70.5\%$ on GSM8K (+3.5 points). Further analysis (Figure 5) for these math reasoning tasks shows that learned summaries promote greater output diversity across different pass@k settings, compared to vanilla MTP with immediate future token prediction. A promising implication of this increased diversity is that post-training on the FSP base model may yield stronger results, as recent work demonstrates that the base model’s solution coverage forms an upper bound on what RL fine-tuning can achieve (Shao et al., 2024; Yue et al., 2025).
@@
In Section˜10.3, we replicate these findings at 3B scale, where both handcrafted and learned future summaries improve over vanilla MTP, again most prominently on math reasoning tasks. Additional ablations explore the effects of omitting reweighting in FSP-BoW and using learned summaries from deeper layers of the ReverseLM teacher model in FSP-RevLM.
(a) 1T Unique Tokens (1 Epoch)
(b) 50B Unique Tokens (20 Epochs)
Figure 6: Data-constrained pretraining at 8B scale. Left: standard pretraining with 1T unique tokens for 1 epoch (same as Table˜1). Right: data-constrained pretraining with 50B unique tokens for 20 epochs, matched for total compute. We report average accuracy across all reasoning tasks. FSP-RevLM outperforms NTP in both regimes, while in the data-constrained setting all future-aware objectives surpass NTP, indicating that richer future-aware supervision is especially useful when unique data are limited.
Data-constrained Pretraining Results.
Motivated by recent work highlighting data-constrained pretraining regimes, in which repeated exposure to a limited pool of unique tokens becomes increasingly common (Muennighoff et al., 2023; Kim et al., 2025), we also evaluate our 8B models in this setting. Specifically, we train on 50B unique tokens for 20 epochs, so that the total token budget matches the 1T-token budget used in the 1-epoch setup considered previously in Table˜1. This enables an iso-compute comparison of pretraining objectives as the supply of unique data is reduced. Figure 6 reports average performance across all reasoning tasks, with the left sub-figure corresponding to the standard 1-epoch regime and the right sub-figure to the data-constrained 20-epoch regime. We find that FSP-RevLM continues to outperform NTP in the data-constrained setting, while remaining competitive with MTP variants. Notably, in the data-constrained regime all future-aware objectives outperform NTP, including DS-MTP, which underperformed in the 1-epoch setting. Overall, these results reinforce the view that when the total unique tokens are limited, richer future-aware supervision can extract more useful learning signal from each token than next-token prediction alone. Further, in Appendix 10.4 we present a more comprehensive analysis at the 3B scale, showing that as the budget of unique tokens decreases, future-aware methods outperform NTP even on standard language modeling metrics.